# Performance Gap in Entity Knowledge Extraction Across Modalities in Vision Language Models

### v1

@@ Results.
Figure 2 shows the effect of patching image tokens across different layers.
@@
Layers 0–19: When image tokens are patched at or before layer 19, the model predominantly predicts the injected entity. This strengthens the hypothesis that the middle layers play a critical role in how the model interprets the image. Our results suggest that despite having 18 layers to process and propagate information about the original entity to later positions, the information conveyed within the narrow window of layers 19 to early twenties ultimately determines the outcome.
Layers 20+: Patching beyond layer 19 results in the model reliably predicting the target entity, suggesting that image-based entity identification is finalized by this stage.
These results suggest two possible processing paths:
@@
I. Parallel identification and knowledge extraction: Generation and query token representations, can be enriched by the image tokens via the attention mechanism concurrently with entity identification before layer 19, implying that the model incrementally combines these processes over several layers, with more impact to information passed in later layers.
@@
II. Sequential processing: Identification may occur just after layer 19, with the remaining layers allocated for factual reasoning and prediction tasks. This would leave a limited subset of late layers for extracting factual knowledge about the entity, potentially leading to degraded question-answering performance.
Figure 2: Entity Patching: Image token hidden states are extracted from the forward pass of the injected entity’s image and injected into the same layer of the original entity’s forward pass. The x axis is the injection layer. Orange line indicates the percent of times the injected entity was generated. Blue is the percent of times the original entity was generated.
3.2 Long processing is necessary in most cases
@@
Having identified the middle layers as the point where critical information propagates from image hidden states to later positions, and having established an upper bound on where identification happens, we will try to answer which of the above mentioned processing paths is more likely, and also answer whether all these layers are truly required for identification? Could it be that the entity was already identified in earlier layers, but was unable to utilize this in the successive layers?
To answer these questions we apply activation patching of the image hidden representations from an early layer into layer 19, by setting $\left(\mathbf{h}_{1,v}^{19}\ldots\mathbf{h}_{n,v}^{19}\right)=\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}\right)$ for values of $\ell$ ranging 0 to 18. If identification indeed happens in early layers, "skipping" all the layers until the point of information propagation should retain most of the identification accuracy.
@@ Results.
Figure 3 shows the results, where every bar shows the percent of correctly identified entities after applying forward patching from the layer indicated by axis x to layer 19. From these results we learn that even when patching the representation from layer 0, 40% of entities are still correctly recognized. This suggests that in these cases the representation provided by the vision encoder, encodes sufficient signals to identify the entity. Up to layer 9 there is very little improvement, indicating that these play a lesser role in the identification process. From layer 10 we see a significant improvement until layer 18 where 80% of the entities are identified. This trend indicates that in the majority of cases, the middle layers (10-19) are critical in the identification process.
With these results in mind, we turn back to our motivating experiment and ask whether the QA accuracy is better for entities with early identification. In this experiment, we split our dataset into entities where forward patching from early $(<10)$ layers maintains correct identification, lets call these early-id, and those requiring late layers, late-id. The average accuracy per entity of the early-id entities is 40% compared to 37% for the late-id. This small difference means that the model does not benefit from early identification and this critical information waits for the middle layers to be attended to. This strengthens the case for the second processing pass suggested before, of sequential processing of identification and extraction.
Figure 3: Each bar shows the percent of correctly identified entities after applying forward patching from the layer indicated by axis x to layer 19.
@@ 4 Related work
VLMs have emerged as a powerful tool, demonstrating impressive capabilities in tasks such as visual reasoning, object recognition, and multimodal understanding Basu et al. (2024). However, their internal mechanisms and interpretability remain active areas of research. Below, we review the related works categorized into general VLM interpretability, cross-modal information flow, attention mechanisms, and mechanistic approaches to visual token processing.
Recent surveys on large multimodal models provide a comprehensive overview of explainability challenges in these systems. Dang et al. (2024) propose a framework categorizing interpretability across data, model, and training perspectives. They highlight the alignment and decomposition of modalities as key challenges, emphasizing the need for explainability to ensure trustworthiness in multimodal AI applications. In this context, techniques for token-level, neuron-level, and embedding-level interpretability have been systematically analyzed, offering insights into the mechanisms driving decision-making processes.
Previous research has examined the internal information distribution within these models Basu et al. (2024), highlighted the limitations arising from pre-trained vision encoders Tong et al. (2024), and investigated the issue of hallucinations in VLMs Bai et al. (2024). Studies such as Huo et al. (2024) delve into neuron-level interpretations, uncovering how specific neurons contribute to multimodal reasoning. Similarly, Jiang et al. (2024) address hallucination issues in VLMs by examining the influence of vision-language representations. Both works highlight the importance of dissecting model components to improve transparency and robustness.
@@ 5 Conclusion
This work explores how VLMs process visual entities, enhancing understanding of multi-modal architectures. Our findings highlight that reasoning about entities shown in an image is not a straightforward extension of reasoning about entities described in text, rather, it more closely resembles the processing of multi-hop questions. The model dedicates a substantial portion of its computational "budget" to entity identification and information processing, leaving few layers for the subsequent factual reasoning step. Thus, critical image-related processing often concludes relatively late in the processing flow, limiting the model’s capacity to extract factual knowledge in the remaining layers thus limiting its overall performance. We also find that this issue persists even for early-id entities, that could be recognized earlier, indicating that current models may not be using their representational capacity efficiently.
By highlighting these challenges, our results offer pathways for improvement. Potential directions for tackling the root of this issue include adjusting training regimes to better separate or balance identification and knowledge extraction steps, as well as designing architectural enhancements that reduce entanglement between these stages, which could lead to VLMs that better, and more reliably, utilize their internal knowledge. For instance, this behavior may be mitigated to some extent by using early-fusion models, such as Chameleon. We leave this exploration for future work.
@@
We focused on the performance of Llava-1.5-7B, as it is a canonical architecture in the field, and allows applying interventions with ease. While Llava is a widely used VLM that has been found useful in many applications, newer, early fusion Team (2024) models may exhibit different behavior to the kind described in this work. We believe that studying the gap we report in this work for Llava also for such early fusion models would be an interesting direction for a future research.
@@
Lastly, while our experiments show clear trends in terms of model performance and internal behaviors, they were conducted on a set of images of 300 popular entities. Extending our analysis to a larger set of entities with different properties, such as popularity and ethnicity, is an important direction for future work.
@@ Acknowledgments
We thank Omri Kaduri for his input on Llava mechanics. IC is supported by the Israeli Science Foundation (grant number 1186/18). DG is supported by the Israeli Science Foundation (grant number 1083/24).
Appendix A Additional experiment: Critical information is processed late
To investigate the stages of the model’s processing pipeline where entity identification occurs, we ask the model to identify the subject in an image and employ an attention knockout strategy following the methodology of Geva et al. (2023). This involves selectively prohibiting attention from textual tokens to image tokens at various layers of the model, effectively blocking cross-modal information flow. We do this by inserting $-inf$ in the attention mask at $[x,y]$ where $x$ are the indices of $\mathbf{h}_{1,q}^{\ell}\dots\mathbf{h}_{m,q}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}$, $y$ are the indices of $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$, and $\ell$ is any layer we want to knock out. We conduct this experiment in two directions:
@@ Appendix A Additional experiment: Critical information is processed late
Results. Figure 4 presents both types of knockout:
@@
Table 1: An example of bottom up knockout on an image of LeBron James. The layer numbers indicate the final knockout layer. Only layers with meaningful change in output were included for brevity. Note that the model briefly recovers accurate entity identification.
Appendix B Additional Qualitative Examples of Attention Knockout

### v2

@@ Results.
Figure 4 shows the effect of patching image hidden states across different layers, indicated by the X axis. The first three graphs show results for patching between entities of the same type, for different models. The three models exhibit different behaviors but with similar trends. The main difference between the models is the layer of the critical point, mentioned earlier (where the blue and orange lines cross). Although for all models this occurs after the middle of the model (layer 16), it varies from 17 for LM-SigLIP to 24 for LM-CLIP. It is important to note the higher standard deviation for the LLaVA-MORE-8B variants. For LLaVA7B the critical point is clearly marked by a sudden change at layer 20, while for LM-CLIP the change from predicting the injected entity to the original is more gradual but still clearly after the middle layer.
@@
In all three cases, when the image tokens are patched at or before the middle layer, the model predominantly predicts the injected entity. This strengthens the hypothesis that the middle layers play a critical role in how the model interprets the image. Our results suggest that despite having 15-20 layers to process and propagate information about the original entity to later positions, the information conveyed within the narrow window of the central layers ultimately determines the outcome. Patching beyond layer 25 results in the model reliably predicting the target entity, suggesting that image-based entity identification is mostly finalized by this stage.
The last (bottom) graph in Figure 4, shows results for patching between different types of entities for LLaVA7B. Note the almost identical behavior to cross patching between same types (top), albeit a slightly faster decline of the orange line. This provides even stronger evidence that the model’s prediction is compelled by the information propagated in latter layers. Even when having 13 layers to attend to the visual hidden states that encode the identity of, say, LeBron James, new information that comes in at layer 14 encoding the logo of Spotify causes the model to predict the latter. Up until layer 14 this happens with a striking  100% of entities, regardless of type.
The results suggest two possible processing paths:
@@
I. Parallel identification and knowledge extraction: Generation and query hidden states, can be enriched by the image hidden states, via the attention mechanism, concurrently with entity identification before the middle layers. This would imply that the model incrementally combines these processes over several layers, with information passed in later layers having more impact.
@@
II. Sequential processing: Identification may occur just after the middle layer, with the remaining layers allocated to factual reasoning and prediction tasks. This would leave a limited subset of late layers for extracting factual knowledge about the entity, potentially leading to degraded question-answering performance.
Figure 4: Results for the experiment in Section 5, illustrated in Figure 2. The top three graphs show results when cross patching between forward passes of entities of the same type, for three different models (indicated in the titles). The bottom graph shows results for cross patching between different types, for LLaVA7B. The X axis indicates the injection layer. The orange line indicates the percent of times the injected entity was generated. The blue line is the percent of times the original entity was generated.
6 Long Processing is Often Necessary
@@
Having identified the middle layers as the point where critical information propagates from image hidden states to later positions, and having established an upper bound on where identification happens, we will try to answer which of the above mentioned processing paths is more likely, and also answer whether all these layers are truly required for identification. Could it be that the entity was identified in earlier layers, but the model was unable to utilize this until later layers?
To answer these questions we apply activation patching of the image hidden states from an early layer into all subsequent layers until layer 20. In detail, at every run we choose a source layer source and set $\left(\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}\right)=\left(\mathbf{h}_{1,v}^{\textit{source}}\ldots\mathbf{h}_{n,v}^{\textit{source}}\right)$ for values of $\ell$ ranging from source to 20. We repeat this for source layers 0 to 19. This effectively "freezes" the hidden states at the image token positions, not allowing them to be updated by neither the attention nor MLP modules. Importantly, this does not prevent the query and generated positions to attend to the image tokens, only that they see the same state at all the patched layers. We choose 20 as the end point since this is past the point we identified in the last experiment, where we claim that entity identification is mostly done and propagated to later positions. If identification indeed happens in early layers, "freezing" all the layers until the point of information propagation should retain most of the identification accuracy.
@@ Results.
Figure 5 shows the results, where every bar shows the percent of correctly identified entities after applying patching from the source layer indicated by the x axis onto all layers until layer 20. There is no distinction here between different types of entities. From these results we see an interesting difference between the models. For LLaVA7B, processing the image hidden states for more layers is required to reach the same results as LM-SigLIP and LM-CLIP do with between 4 and 10 layers less. All layers seem to take some part in this process, as every layer allows identifying more entities. As expected, this gap diminishes the more layers are allowed for processing. When patching through layers 13 to 20, LLaVA7B catches up to the other models. Interestingly, for both LLaVA-MORE-8B variants, patching through all layers from 0 until 20, still retains almost 50% (LM-SigLIP) and over 30% (LM-CLIP) of recognized entities. This suggests that in these cases, the representation provided by the visual backbone encodes a sufficient signal to identify the entity. This alone would not be enough without the LLM being able to utilize this representation, suggesting better alignment of the visual and textual embedding spaces. These results complement those of the previous experiment showing that LM-SigLIP, which had the earliest critical point in the previous experiments, can identify the most entities with the least amount of processing (manifested in number of layers).
With these results in mind, we turn back to our motivating experiment and ask whether early identification leads to better QA accuracy. As the graph shows, this question is mostly relevant for the LLaVA-MORE-8B variants. In this experiment, we split our dataset into entities where forward patching from early $(<5)$ layers maintains correct identification. We refer to these as early-id, and those requiring late layers, late-id.
Table 2 shows the results for both models on both splits. Notice that the early-id entities achieve higher accuracy in both models and this rise is consistent for both the image and textual representation. This consistency between modalities teaches us that better knowledge of an entity (higher accuracy) leads to early recognition, and not the other way around. Had that been the case, we would expect to see a rise in image accuracy for early-id entities without a corresponding rise in text accuracy. This leads to the accuracy drop between image accuracy and text accuracy to actually be higher for the early-id entities. This means that although the model has access to the entity’s identity at an early layer, leaving more layers for the second hop, the model does not utilize this advantage. This could indicate that although this information exists, it does not reach the query tokens via the attention mechanism early enough to inform subsequent factual recall. This strengthens the observation in Section 5 that information propagates in the middle layers.
Figure 5: Results for the experiment in Section 6, illustrated in Figure 2. Each bar shows the percent of correctly identified entities after applying patching from the layer indicated by the x axis into all layers up to layer 19.
Model Split Img Acc Txt Acc Drop
LM-CLIP Early 0.260 0.442 0.182
LM-CLIP Late 0.192 0.347 0.154
LM-SigLIP Early 0.234 0.417 0.183
LM-SigLIP Late 0.168 0.338 0.170
Table 2: Results on POPVQA, split by early and late identified entities. Early identification is correlated with higher accuracy but not with lower drop.
@@ 7 Related Work
While VLMs demonstrate impressive capabilities in tasks such as visual reasoning, object recognition, and multimodal understanding Basu et al. (2024), their internal mechanisms and interpretability remain active areas of research. Below, we review the related works categorized into general VLM interpretability, cross-modal information flow, attention mechanisms, and mechanistic approaches to visual token processing.
Dang et al. (2024) survey VLMs proposing a framework categorizing interpretability across data, model, and training perspectives. They highlight the alignment and decomposition of modalities as key challenges, emphasizing the need for explainability to ensure trustworthiness in multimodal AI applications. In this context, techniques for token-level, neuron-level, and embedding-level interpretability have been systematically analyzed, offering insights into the mechanisms driving decision-making processes.
Previous research has examined the internal information distribution within these models Basu et al. (2024), highlighted the limitations arising from pre-trained vision encoders Tong et al. (2024), and investigated the issue of hallucinations in VLMs Bai et al. (2024); Jiang et al. (2025). Studies such as Huo et al. (2024) delve into neuron-level interpretations, uncovering how specific neurons contribute to multimodal reasoning. Similarly, Jiang et al. (2024) address hallucination issues in VLMs by examining the influence of vision-language representations. Both works highlight the importance of dissecting model components to improve transparency and robustness.
@@
Some datasets probe how well VLMs handle knowledge-intensive queries Lerner et al. (2024); Zhu et al. (2024); Onoe et al. (2025); Yanuka et al. (2025); Wei et al. (2025). For instance, InfoSeek Chen et al. (2023) is a VQA dataset tailored for information-seeking questions that cannot be resolved using only common sense knowledge. Their findings reveal that state-of-the-art VLMs often struggle with these questions; however, fine-tuning on InfoSeek enables them to leverage more fine-grained knowledge learned during pre-training. Unlike InfoSeek, which focuses on direct information-seeking queries, our dataset explicitly includes a bridge entity as part of a two-hop reasoning process—requiring both identification of the entity and extraction of the relevant fact. This design allows us to disentangle the efficacy of visual recognition from the subsequent factual retrieval and more directly analyze the gap between textual and visual representations.
@@
Parcalabescu and Frank (2023) introduce a method that uses Shapley values to quantify how VLMs balance information from each modality, revealing varying degrees of unimodal collapse. Parcalabescu and Frank (2025) examine how VLM decoders rely on visual and textual inputs differently when producing answers versus explanations, noting greater reliance on text overall but an increased role for images in explanation tasks. Unlike these approaches, which focus on measuring how images and text are combined in general, our work studies a specific two-step reasoning process—entity identification followed by factual retrieval—and highlights a distinct performance gap when relying on visual instead of textual cues. Zhu et al. (2024) propose a pipeline to detect and mitigate what they refer to as parametric knowledge conflicts in large vision-language models, introducing a dynamic contrastive decoding method that improves performance on InfoSeek. While their work addresses cross-modal inconsistencies by focusing on parametric conflicts, we contribute a distinct perspective centered on bridging entity identification and factual retrieval.
Interestingly, when compared to human behavior, the behavior of VLMs we find in this paper contrasts with findings from cognitive psychology. Specifically, humans tend to perform better with visual rather than textual information, a phenomenon known as the picture superiority effect (Jenkins et al., 1967). While explanations for this effect are still debated, the dual-coding theory (Paivio and Csapo, 1973) attributes this to images being encoded both visually and verbally, unlike text which is only encoded verbally. These findings suggest that humans benefit more from visual input, in contrast to VLMs, which are likely biased toward text due to its prevalence in their training data. We leave a further exploration of the relation between the two to future research.
@@ 8 Conclusion
This work explores how VLMs process visual entities, enhancing understanding of multi-modal architectures. Our findings highlight that reasoning about entities shown in an image is not a straightforward extension of reasoning about entities described in text, rather, it more closely resembles the processing of multi-hop questions. Across different model and visual backbones we see that models dedicate a substantial portion of their computational "budget" to entity identification and information processing, leaving few layers for the subsequent factual reasoning step. Thus, critical image-related processing often concludes relatively late in the processing flow, limiting the model’s capacity to extract factual knowledge in the remaining layers thus limiting its overall performance. We also find that this issue persists even for entities that could be identified earlier, indicating that current models may not be using their representational capacity efficiently.
By highlighting these challenges, our results offer pathways for improvement. Potential directions for tackling the root of this issue include adjusting training regimes to better separate or balance identification and knowledge extraction steps, as well as designing architectural enhancements that reduce entanglement between these stages, which could lead to VLMs that better, and more reliably, utilize their internal knowledge. For instance, this behavior may be mitigated to some extent by using early-fusion models, that aim to create a shared embedding space between the visual and textual modalities. We leave this exploration for future work.
@@
In the second part of our experiments, we focused on the performance of LLaVA style models, as it is a canonical architecture in the field, and allows applying interventions with ease. While we did observe a gap in other architectures, we leave the probing experimentation on these for future work. Indeed, we did not test all type of VLMs, where some may exhibit different behavior to the kind described in this work. We believe that studying the gap we report in this work for a larger variety of VLM types would be an interesting direction for a future research.
@@
Lastly, extending our analysis to entities with different properties and of additional ethnicities is an important direction for future work.
@@ Acknowledgments
We thank Omri Kaduri and Moran Yanuka for their input and feedback. This work was supported by the Tel Aviv University Center for AI and Data Science (TAD) and the Israeli Science Foundation.
Appendix A Motivating Experiment - Results Breakdown By Type
Tables 3 and 4 show a breakdown of the results in section 4 according to type. The first shows the image and text accuracy and drop, while the second shows a fine grained p-value analysis for the Wilcoxon signed-rank test. This breakdown provides some interesting insights. The drop in the paintings case is often low, with Qwen2-VL even doing better with the visual representation. This seems to be due to the model being able to provide more informed guesses for some of the questions. Other than that, there is no correlation between specific types and accuracy drop. It seems that different models have different strengths and weaknesses, for example LM-SigLIP recognizes significantly less brand logos despite the name of the brand often appearing on the logo.
Type #Identified Img Acc Txt Acc Drop
LLaVA7B
Brands 343 0.270 0.432 0.162
Celebs 294 0.342 0.543 0.201
Landmarks 168 0.280 0.502 0.222
Paintings 120 0.128 0.299 0.171
LM-CLIP
Brands 364 0.238 0.374 0.136
Celebs 128 0.278 0.379 0.101
Landmarks 212 0.250 0.486 0.236
Paintings 140 0.083 0.249 0.165
LM-SigLIP
Brands 237 0.169 0.406 0.237
Celebs 136 0.301 0.454 0.153
Landmarks 149 0.259 0.459 0.200
Paintings 138 0.111 0.207 0.096
LLaVA34B
Brands 416 0.546 0.630 0.083
Celebs 415 0.564 0.742 0.179
Landmarks 346 0.534 0.655 0.122
Paintings 109 0.435 0.455 0.020
Qwen2-VL
Brands 462 0.363 0.411 0.049
Celebs 1240 0.418 0.544 0.125
Landmarks 1059 0.474 0.489 0.015
Paintings 382 0.445 0.276 -0.168
Table 3: Fine-grained results on POPVQA broken down by different entity types. Reported accuracy is first calculated per subject then averaged across the entire dataset.
Split LLaVA7B LM-CLIP LM-SigLIP
Overall $3.83\times 10^{-76}$ $1.42\times 10^{-49}$ $6.36\times 10^{-47}$
Brands $1.07\times 10^{-21}$ $1.16\times 10^{-16}$ $4.78\times 10^{-21}$
Celebs $3.48\times 10^{-40}$ $1.34\times 10^{-8}$ $6.87\times 10^{-12}$
Landmarks $4.88\times 10^{-18}$ $6.22\times 10^{-22}$ $2.90\times 10^{-13}$
Paintings $2.51\times 10^{-7}$ $1.29\times 10^{-8}$ $7.33\times 10^{-5}$
Split LLaVA-34B Qwen2-VL
Overall $4.27\times 10^{-56}$ $4.02\times 10^{-29}$
Brands $9.92\times 10^{-11}$ $7.60\times 10^{-5}$
Celebs $3.42\times 10^{-50}$ $1.33\times 10^{-115}$
Landmarks $3.59\times 10^{-13}$ $1.17\times 10^{-3}$
Paintings $3.11\times 10^{-1}$ $1.00$
Table 4: Fine-grained p-value analysis of Wilcoxon signed-rank test between the per-entity accuracies for visual versus textual representations, broken down by different entity types.
Appendix B Additional Experiment: Critical Information is Propagated Late
To investigate the stages of the model’s processing pipeline where entity identification occurs, we ask the model to identify the subject in an image and employ an attention knockout strategy following the methodology of Geva et al. (2023). This involves selectively prohibiting attention from textual tokens to image tokens at various layers of the model, effectively blocking cross-modal information flow. We do this by inserting $-inf$ in the attention mask at $[x,y]$ where $x$ are the indices of $\mathbf{h}_{1,t}^{\ell}\dots\mathbf{h}_{m,t}^{\ell},\mathbf{h}_{1,g}^{\ell}\dots\mathbf{h}_{k,g}^{\ell}$, $y$ are the indices of $\mathbf{h}_{1,v}^{\ell}\ldots\mathbf{h}_{n,v}^{\ell}$, and $\ell$ is any layer we want to knock out. We conduct this experiment in two directions:
@@ Appendix B Additional Experiment: Critical Information is Propagated Late
Results. Figure 6 presents both types of knockout on LLaVA7B:
@@
Table 5: An example of bottom-up knockout on an image of LeBron James. The layer numbers indicate the final knockout layer. Only layers with meaningful change in output were included for brevity. Note that the model briefly recovers accurate entity identification.
Appendix C Additional Qualitative Examples of Attention Knockout
@@ Appendix D Resources
Appendix E Ethics of Data Collection
Note that the data was collected with only permissive licenses. Moreover, we have an approval from our institution IRB for "Checking and understanding the performance of deep neural networks".